# Convexity on manifolds without focal points and spectrum of the Laplacian

### v1

@@ 2. Preliminaries
As $\nabla b_{v}$ is a $C^{1}$ vector field with $\left\lVert\nabla b_{v}\right\rVert=1$, we have that $\nabla_{\nabla b_{v}}\nabla b_{v}=0$. Indeed ${\langle\nabla_{X}\nabla b_{v}},{Y\rangle}={\mathrm{H}}^{b_{v}}(X,Y)$ is symmetric in $X,Y$ and hence for every vector field $X$
@@
In particular, we see that every integral curve of $\nabla b_{v}$ is a geodesic and hence complete. Then, the orthogonal decomposition $T_{p}M={\mathbb{R}}\nabla b_{v}\oplus(\nabla b_{v})^{\perp}$ enables us to write $x={\langle x},{\nabla b_{v}(p)\rangle}_{T_{p}M}\nabla b_{v}(p)+x^{\perp}$. Using $\nabla_{\nabla b_{v}}\nabla b_{v}=0$ and the bilinearity of ${\mathrm{H}}^{b_{v}}$ we have
${\mathrm{H}}^{b_{v}}(p)(x,x)={\mathrm{H}}^{b_{v}}(p)(x^{\perp},x^{\perp}).$
@@
Notice in the above computation that when $X=\nabla b_{v}$ we have ${\mathrm{H}}^{b_{v}}(\nabla b_{v},\nabla b_{v})=0$ and hence $b_{v}$ is not strictly convex for any $v\in SM$. In fact, to the best of our knowledge, it is not known if Busemann functions are even convex for an arbitrary manifold of continuous asymptote. Using Lemma 2.1 and [13, Proposition 2, Theorem 1] we have the following relation between the stable Jacobi tensors and the Busemann functions in the case of manifolds of continuous asymptote: for $v\in SM$, $p\in M$, and $x\in T_{p}M$,
@@
(2.2) ${\mathrm{H}}^{b_{v}}(p)(x,x)=-{\langle D_{\nabla b_{v}(p)}^{{}^{\prime}}(0)x^{\perp}},{D_{\nabla b_{v}(p)}(0)x^{\perp}\rangle}$
Therefore for asymptotically harmonic manifolds, by taking traces in (2.2), we have $\Delta b_{v}=h=trace(D_{v}^{{}^{\prime}}(0))$ for every $v\in SM$.
@@
3. A strictly convex function on a manifold without focal points
A complete simply connected Riemannian manifold $M$ is said to have no focal points if for every nontrivial Jacobi field $Y$ satisfying $Y(0)=0$, the map $t\mapsto\left\lVert Y(t)\right\rVert$ is strictly increasing for $t>0$ [25, Proposition 4]. In particular, any non-trivial Jacobi field cannot vanish more than once and hence every manifold without focal points is also a manifold without conjugate points. Note that $t\mapsto\left\lVert D_{v}(t)\right\rVert$ is monotonically decreasing for each $v\in SM$. This in particular shows that $M$ is of continuous asymptote [13, §5].
Henceforth in this section we assume that $M$ is a $n$-dimensional complete simply connected Riemannian manifold without focal points. To begin with let us recall the geodesic compactifications of $M$ and its connections to the Riemannian exponential map and the Busemann functions. We refer to [16] for more details on what follows. Given $v\in SM$ there exists a unique geodesic $\gamma_{v}$ in $M$ satisfying $\gamma_{v}^{{}^{\prime}}(0)=v$. Two vectors $v,w\in SM$ are said to be asymptotic if the distance between $\gamma_{v}(t)$ and $\gamma_{w}(t)$ is bounded for all $t\geq 0$. Let $M(\infty)$ be the set of all classes of asymptotic vectors and let $\overline{M}=M\cup M(\infty)$. $\overline{M}$ has a canonical topology with the following property: for any $p\in M$, the exponential map ${\mathrm{Exp}}_{p}:T_{p}M\to M$ extends uniquely to a homeomorphism $\Phi_{p}:\overline{T_{p}M}\to\overline{M}$ where $\overline{T_{p}M}\backslash T_{p}M$ is homeomorphic with the unit sphere $S_{p}M$ in $T_{p}M$. Then for any $p\in M$, there exists a unique geodesic $\gamma_{{}_{\nabla b_{v}(p)}}$ starting at $p$ asymptotic to $\gamma_{v}$. From our discussions in §2, we see that the Busemann boundary $\partial^{o}M$ and the geodesic compactification $M(\infty)$ are homeomorphic.
Further, for each $v\in SM$, the Busemann function $b_{v}$ is convex [13, Theorem 2] i.e., ${\mathrm{H}}^{b_{v}}$ is nonnegative definite for every $v\in SM$. Hence each horosphere is a convex submanifold in $M$. As remarked in the introduction, motivated by the construction on symmetric spaces of noncompact type of a strictly convex function, indicated in [10, penultimate paragraph p. 657], we now construct a strictly convex function out of the Busemann functions, using the geometry at infinity for manifolds without focal points discussed above.
@@ 3. A strictly convex function on a manifold without focal points
Let $M$ be a complete non-compact simply connected Riemannian manifold with no focal points and negative Ricci curvature. Then $M$ admits a $C^{2}$-strictly convex function $F$ whose gradient is bounded.
@@
Let $o$ be an arbitrary but fixed point in $M$. Let $\{v_{k}\}$ be a countable dense subset in $S_{o}M$. Consider the Dirac measures $\delta_{v_{k}}$ associated to each $v_{k}$. Then ${\eta=\sum\limits_{k=1}^{\infty}\frac{1}{2^{k}}\delta_{v_{k}}}$ defines a probability measure on $S_{o}M$. We identify $S_{o}M$, $M(\infty)$ and $\partial^{o}M$ via the respective homeomorphisms. Let $F(p)=\displaystyle{\int\limits_{M(\infty)}b_{v}(p)~{}\mathrm{d}\eta(v)}$. By dominated convergence theorem, the gradient of $F$ at $p$ is given by $\nabla F(p)=\displaystyle{\int\limits_{M(\infty)}\nabla b_{v}(p)~{}\mathrm{d}\eta(v)}$. Clearly,
$\left\lVert\nabla F(p)\right\rVert=\left\lVert\int\limits_{M(\infty)}\nabla b_{v}(p)~{}\mathrm{d}\eta(v)\right\rVert\leq\int\limits_{M(\infty)}\left\lVert\nabla b_{v}(p)\right\rVert~{}\mathrm{d}\eta(v)=1,\quad\text{as}~{}\left\lVert\nabla b_{v}\right\rVert=1$
i.e., $F$ has bounded gradient. In particular, $F$ is Lipschitz with Lipschitz constant $1$. For $p\in M$ and $w\in S_{p}M$, using (2.2) and that ${\mathrm{H}}^{b_{v}}$ is nonnegative definite we have
$\int\limits_{M(\infty)}{\mathrm{H}}^{b_{v}}(p)(w,w)~{}\mathrm{d}\eta(v)$ $=\left|\int\limits_{M(\infty)}{\langle D_{\nabla b_{v}(p)}^{{}^{\prime}}(0)w},{w\rangle}~{}\mathrm{d}\eta(v)\right|$
$\leq\int\limits_{M(\infty)}|{\langle D_{\nabla b_{v}(p)}^{{}^{\prime}}(0)w},{w\rangle}|~{}\mathrm{d}\eta(v)$

### v2

@@ 2. Preliminaries
As $\nabla b_{v}$ is a $C^{1}$ vector field with $\left\lVert\nabla b_{v}\right\rVert=1$, we have that $\nabla_{\nabla b_{v}}\nabla b_{v}=0$. Indeed ${\langle\nabla_{X}\nabla b_{v}},{Y\rangle}=\mathrm{Hess}\penalty 10000\ b_{v}(X,Y)$ is symmetric in $X,Y$ and hence for every vector field $X$
@@
In particular, we see that every integral curve of $\nabla b_{v}$ is a geodesic and hence complete. Then, the orthogonal decomposition $T_{p}M={\mathbb{R}}\nabla b_{v}\oplus(\nabla b_{v})^{\perp}$ enables us to write $x={\langle x},{\nabla b_{v}(p)\rangle}_{T_{p}M}\nabla b_{v}(p)+x^{\perp}$. Using $\nabla_{\nabla b_{v}}\nabla b_{v}=0$ and the bilinearity of $\mathrm{Hess}\penalty 10000\ b_{v}$ we have
$\mathrm{Hess}\penalty 10000\ b_{v}(p)(x,x)=\mathrm{Hess}\penalty 10000\ b_{v}(p)(x^{\perp},x^{\perp}).$
@@
Notice in the above computation that when $X=\nabla b_{v}$ we have $\mathrm{Hess}\penalty 10000\ b_{v}(\nabla b_{v},\nabla b_{v})=0$ and hence $b_{v}$ is not strictly convex for any $v\in SM$. In fact, to the best of our knowledge, it is not known if Busemann functions are even convex for an arbitrary manifold of continuous asymptote. Using Lemma 2.1 and [Esc77, Proposition 2, Theorem 1] we have the following relation between the stable Jacobi tensors and the Busemann functions in the case of manifolds of continuous asymptote: for $v\in SM$, $p\in M$, and $x\in T_{p}M$,
@@
(3) $\mathrm{Hess}\penalty 10000\ {b_{v}}(p)(x,x)=-{\langle D_{\nabla b_{v}(p)}^{{}^{\prime}}(0)x^{\perp}},{D_{\nabla b_{v}(p)}(0)x^{\perp}\rangle}$
In what follows, we shall make use of the foregoing discussion for simply connected Riemannian manifolds without focal points and henceforth make this assumption.
3. A construction using the distance function and its applications
In this section, we consider the distance from a fixed point on a simply connected manifold $M$ without focal points, and the strictly convex functions derived from it, having additional properties required for our applications. For a fixed $p\in M$, the function $q\mapsto d^{2}(p,q)$ is smooth and strictly convex where $r(q)=d(p,q)$ is the distance function. This is proved by an analysis of Jacobi fields as follows. Note that the normalized distance-squared function $u(q)=\frac{1}{2}r^{2}(q)$ is smooth on $M$. The Hessian of $u$ is given by
$\mathrm{Hess}\penalty 10000\ u=\mathrm{d}r^{2}+r\mathrm{Hess}\penalty 10000\ r$
Given $v\in T_{p}M$, consider the Jacobi field along $t\mapsto\exp_{p}(tv)$ satisfying $Y(0)=0$. Then a routine calculation yields,
(4) $\mathrm{Hess}\penalty 10000\ r(\exp_{p}(tv))(Y(t),Y(t))={\langle Y^{{}^{\prime}}(t)},{Y(t)\rangle}=\frac{1}{2}\frac{d}{ds}\Big|_{s=t}\left\lVert Y(s)\right\rVert^{2}$
Since $M$ has no focal points, $d\exp_{p}$ is nonsingular at $tv$ for any $w\in T_{q}M$ there exists a Jacobi field $Y$ along $\gamma$ such that $Y(0)=0$ and $Y(t)=w$. Using [O’S74, Theorem 4] we see that $\mathrm{Hess}\penalty 10000\ r>0$ for $t>0$ and so $u$ is strictly convex on $M$ and we refer to [Ebe72] for more details. It then follows from [Yam82, Theorem A] that each compact subgroup of the isometry group of $M$ has a common fixed point. This generalizes the classical theorem of Cartan for Hadamard manifolds.
Now we construct, in the usual way, a smooth strictly convex function on $M$ whose gradient is bounded by taking $f(q)=\sqrt{1+2u(q)}$; then $\left\lVert\nabla f\right\rVert\leq 1$ and
$\mathrm{Hess}\penalty 10000\ f$ $=\frac{1}{(1+2u)^{\frac{3}{2}}}\mathrm{d}r^{2}+\sqrt{\frac{2u}{1+2u}}\mathrm{Hess}\penalty 10000\ r$
is positive definite as $\mathrm{Hess}\penalty 10000\ r$ is strictly positive showing that $f$ is strictly convex. A real-valued function $\varphi$ on a manifold is called an exhaustion function if all its sub-level sets $\{\varphi\leq c\}$ are compact. Notice that the sub-level sets of $f$ are given by
$f^{-1}(c)=\begin{cases}\emptyset,&\text{if}\penalty 10000\ c<1\\ \{p\},&\text{if}\penalty 10000\ c=1\\ B(p,c),&\text{if}\penalty 10000\ c>1\end{cases}$
and hence are all compact. Here $B(p,c)$ denotes the closed ball of radius $c$ centred at $p$. Thus $f$ is also an exhaustion function on $M$. For convenience, we summarize this discussion as a lemma.
Lemma 3.1.
@@ 3. A construction using the distance function and its applications
Let $M$ be a simply connected Riemannian manifold without focal points. Then it possesses a smooth strictly convex exhaustion function with bounded gradient.
As a first application we have the following result on the isoperimetric profile function $\mathcal{I}_{M}$ on $M$. The isoperimetric profile of a Riemannian manifold is the function that assigns to a given positive volume the infimum of the perimeter of the sets of this volume. We remark here that regularity properties of this function are of interest and refer to the introduction in [Rit17] for a detailed discussion. The continuity and the nondecreasing nature of the isoperimetric profile is shown [Rit17, Theorem 3.1] for complete manifolds which possess smooth strictly convex Lipschitz exhaustion functions. It then follows from a theorem of Lebesgue that $\mathcal{I}_{M}$ is differentiable almost everywhere on such manifolds. As examples there, Hadamard manifolds and manifolds with strictly positive sectional curvatures are discussed. We give below a class of examples for which the continuity of $\mathcal{I}_{M}$ holds while making no assumptions on curvature and which properly contains Hadamard manifolds. As noted before this class contains plenty of examples whose sectional curvatures do not have a fixed sign [Gul75].
Proposition 3.2.
Simply connected Riemannian manifolds without focal points have a continuous nondecreasing isoperimetric profile function. In particular, they are differentiable almost everywhere.
Proof.
By Lemma 3.1, $M$ admits a smooth strictly convex Lipschitz exhaustion function. Invoking [Rit17, Theorem 3.1], we see that $\mathcal{I}_{M}$ is nondecreasing and continuous. ∎
Regularity for $\mathcal{I}_{M}$ has many interesting consequences; for instance, we have the following result whose proof makes use of standard arguments.
Proposition 3.3.
Let $M$ be as in Proposition 3.2. Suppose $v_{0}$ is a differentiable point of $\mathcal{I}_{M}$. If $\Omega$ is an isoperimetric region in $M$ of volume $v_{0}$, then (the regular part of) its boundary $\partial\Omega$ has constant mean curvature $H=\mathcal{I}_{M}^{{}^{\prime}}(v_{0})$.
Proof.
Since $\Omega$ is isoperimetric region the (regular part of) $\partial\Omega$ has constant mean curvature $H$. Let $\mu,\nu$ be the outward and inward normal variational vector fields on $\partial\Omega$ with compact support. Then the first variations of the volumes and the boundary areas are as follows
$\frac{d}{dt}V(\Omega_{t})\Big|_{t=0}$ $=\int\limits_{\partial\Omega_{0}}|\mu|\mathrm{d}A,\quad\frac{d}{dt}V(\Omega_{t})\Big|_{t=0}=-\int\limits_{\partial\Omega_{0}}|\nu|\mathrm{d}A$
$\frac{d}{dt}A(\partial\Omega_{t})\Big|_{t=0}$ $=H\int\limits_{\partial\Omega_{0}}|\mu|\mathrm{d}A,\quad\frac{d}{dt}A(\partial\Omega_{t})\Big|_{t=0}=-H\int\limits_{\partial\Omega_{0}}|\nu|\mathrm{d}A$
Hence, using the differentiability assumption at $v_{0}$,
$\lim\limits_{\delta v\to 0^{+}}\frac{\mathcal{I}_{M}(v_{0}+\delta v)-\mathcal{I}_{M}(v_{0})}{\delta v}\leq H,\quad\lim\limits_{\delta v\to 0^{-}}\frac{\mathcal{I}_{M}(v_{0}+\delta v)-\mathcal{I}_{M}(v_{0})}{\delta v}\geq H$
yields $\mathcal{I}_{M}^{{}^{\prime}}(v_{0})=H$. ∎
As a consequence, all isoperimetric regions in $M$ of the same volume $v_{0}$ must also have the same mean curvature if $v_{0}$ is a differentiable point of the isoperimetric profile function. In general, the existence of isoperimetric domains of a given volume in a complete noncompact Riemannian manifold is a difficult problem (see [Rit23]).
We now discuss a couple of direct applications in the situation when $M$ is, in addition, Kähler. For general facts on Kähler manifolds we refer to [GW77, §7]. On a Kähler manifold the complex analogue of convex functions is given by the notion of plurisubharmonic functions. These are defined using the complex Hessian called the Levi form. In local coordinates, the Levi form of a function $\tilde{f}$ is given by
$L\tilde{f}=4\sum\limits_{i,j=1}^{n}\frac{\partial^{2}h}{\partial z_{i}\partial\overline{z}_{j}}\mathrm{d}z_{i}\otimes\mathrm{d}\overline{z}_{j}$
where $(z_{1},\cdots,z_{n})$ is a local holomorphic coordinate system. Let $(x_{1},y_{1},\cdots,x_{n},y_{n})$ be the corresponding normal coordinates with $z_{k}=x_{k}+iy_{k}$, $k=1,\cdots,n$. A real-valued function on a Kähler manifold is said to be (strictly) plurisubharmonic, if the Levi form is positive (definite) semi-definite. Since $M$ is Kähler, the Levi form $L_{\tilde{f}}$ of $\tilde{f}$ is expressed in terms of Hessian of $\tilde{f}$ using the real coordinates as
$L_{\tilde{f}}=\mathrm{Hess}(\tilde{f})\left(\frac{\partial}{\partial x_{1}},\frac{\partial}{\partial x_{1}}\right)+\mathrm{Hess}(\tilde{f})\left(\frac{\partial}{\partial y_{1}},\frac{\partial}{\partial y_{1}}\right).$
Therefore if $\tilde{f}$ is a strictly convex $C^{2}$-function on a Kähler manifold, then it is strictly plurisubharmonic.
Stein manifolds are complex manifolds which are biholomorphic to a closed complex submanifold of $\mathbb{C}^{n}$ for some $n\in\mathbb{N}$. This tells us that there are plenty of holomorphic functions on such manifolds. Stein manifolds are necessarily Kähler and a question of great interest then is to ask which Kähler manifolds are Stein. Typically, sufficient conditions are given by putting constraints on the sign of of sectional, holomorphic bisectional or Ricci curvatures[GW77, Theorem $1$]. It is in this context that we observe the following sufficient condition where we make no assumptions on curvature.
Proposition 3.4.
Let $M$ be a simply connected Riemannian manifold without focal points which is also Kähler. Then $M$ is necessarily Stein.
Proof.
Lemma 3.1 gives a smooth strictly convex exhaustion function on $M$, while the discussion above on the Levi form shows that this function is a strictly plurisubharmonic exhaustion function on $M$. Then Grauert’s proof of Levi’s problem [GW77, §4] that a complex manifold is Stein if and only if it admits a strictly plurisubharmonic exhaustion function concludes the proof. ∎
As another application of Lemma 3.1, we have the following volume growth comparison for metric balls in Kähler manifolds without focal points using [NT03, Lemma 4.2 (i)].
Proposition 3.5.
Let $M$ be a simply connected Riemannian manifold without focal points which is also Kähler and of (complex) dimension $m$. Then $M$ admits a smooth strictly plurisubharmonic function with bounded gradient. Moreover, for each $p\in M$ there exists a constant $C_{2}>0$ such that,
$V_{p}(r)\geq C_{2}r^{m},\quad r\geq 1$
where $V_{p}(r)$ is the volume of a metric ball of radius $r$ about $p$ in $M$.
As mentioned in the introduction, the question of whether a simply connected harmonic manifold that carries a Kähler structure must be a symmetric space is still unresolved. However, in the context of Damek–Ricci spaces this has been verified. Therefore, at least in certain special cases such as harmonic manifolds without focal points, the resulting volume growth reflects the same rigidity, in accordance with Bishop–Günther’s comparison theorem [Cha06, Theorem III.4.2].
Our next application is spectral theoretic in nature, and so we begin with a sketch of the basic functional analytic setup concerning the spectrum of the Laplacian referring to [Bor20] for the details. Let $M$ be a complete Riemannian manifold. Then the (positive) Laplacian $\Delta$ is an essentially self-adjoint operator with the space of compactly supported smooth functions on $M$ as the domain. We also denote its unique self-adjoint extension, having the domain $\mathcal{D}(\Delta):=\{f\in L^{2}(M):\Delta f\in L^{2}(M)\}$, by $\Delta$. Using the functional calculus for $\Delta$ we have the family of spectral projection operators $\prod_{E}:=\chi_{E}(\Delta)$ for each Borel subset $E$ of $\mathbb{R}$} referred to as the resolution of $\Delta$. The support of these projections and the spectrum of $\Delta$ are related in the following way. We say that $\lambda\in{\mathbb{C}}$ is in the spectrum $\sigma(\Delta)$ of $\Delta$ if and only if $\prod_{(\lambda-\epsilon,\lambda+\epsilon)}\neq 0$ for all $\epsilon>0$. If $\prod_{\{\lambda\}}\neq 0$, then $\lambda$ is an eigenvalue of $\Delta$ and the range of $\prod_{\{\lambda\}}$ is the corresponding eigenspace. The spectral theorem gives a unitary equivalence between $\Delta$ and a multiplication operator on a suitable Hilbert space of the form $\mathrm{L}^{2}(X,\mu)$, where $\mu$ is a $\sigma$- finite Borel measure on a measure space $X$. By the Lebesgue decomposition, $\mu$ decomposes as $\mu=\mu_{pp}+\mu_{ac}+\mu_{sc}$ into pure point measure, absolutely continuous and singularly continuous measures with respect to the Lebesgue measure, and consequently we have that
(5) $\mathrm{L}^{2}(X,\mu)=\mathrm{L}^{2}(X,\mu_{pp})\oplus\mathrm{L}^{2}(X,\mu_{ac})\oplus\mathrm{L}^{2}(X,\mu_{sc}).$
Here we take the closure for the set of all eigenvalues as it is not necessarily closed. We refer to $\sigma_{ac}(\Delta)$ and $\sigma_{sc}(\Delta)$ as the absolutely continuous and singularly continuous parts of $\sigma(\Delta)$ respectively and their union as the continuous part $\sigma_{cont}(\Delta)$ of $\sigma(\Delta)$. There is also a decomposition of $\sigma(\Delta)$ as a disjoint union:
$\sigma(\Delta)=\sigma_{disc}(\Delta)\cup\sigma_{ess}(\Delta)$
where $\sigma_{disc}(\Delta)$ and $\sigma_{ess}(\Delta)$ denote the discrete and essential parts of $\sigma(\Delta)$. The former consists of all eigenvalues of finite multiplicities and hence is contained in $\sigma_{pt}(\Delta)$. The latter is the complement of $\sigma_{disc}(\Delta)$ in $\sigma(\Delta)$ thereby we have that the continuous part of the spectrum, namely $\sigma_{ac}(\Delta)\cup\sigma_{sc}(\Delta)$ is contained in $\sigma_{ess}(\Delta)$. Note that it is possible that $\sigma_{ess}(\Delta)$ contains eigenvalues of infinite multiplicities usually referred to as the embedded eigenvalues.
In order to analyse the nature of the spectrum of the Laplacian using convexity, we make use of [Xav88, Theorem 3]. The idea there was to use Kato’s $H$-smooth operators to show that if $M$ admits a strictly convex function, then certain boundedness conditions on its gradient and Laplacian implies the non-triviality of the absolutely continuous part of the spectrum. This is then applied in the context of Hadamard manifolds with certain additional assumptions which are satisfied with lower bounds on sectional curvatures, to show the nontriviality of the absolutely continuous spectrum. It is this result that we now generalise to simply connected Riemannian manifolds without focal points having nonpositive radial curvature with respect to a single point. This is a much weaker condition than the nonpositivity of sectional curvature. To this end, we prove the following lemma.
Lemma 3.6.
Let $M$ be a simply connected Riemannian manifold without focal points. Suppose there exists a point $p$ in $M$ such that the radial curvatures with respect to $p$ are nonpositive. Then we have for any $q\neq p$ in $M$
$r(q)\mathrm{Hess}\penalty 10000\ r(q)(v,v)\geq\left\lVert v\right\rVert^{2}\qquad\forall\penalty 10000\ v\in T_{q}M$
Proof.
Since $M$ has no focal points $\exp_{p}:T_{p}M\to M$ is a diffeomorphism. Hence for any $q\in M$, $q\neq p$, there exists a unique $t\in{\mathbb{R}}$ and a unit tangent vector $v\in T_{p}M$ such that $q=\exp_{p}(tv)$. Moreover, $\gamma(s)=\exp_{p}(sv)$ is a unit speed geodesic in $M$ joining $p$ and $q$ satisfying $\gamma(0)=p$ and $\gamma(t)=q$. Then $\gamma^{\prime}(s)=\nabla r(\gamma(s))$ and in particular, $\gamma^{\prime}(t)=\nabla r(q)$. Since $d\exp_{p}$ is nonsingular at $tv$, for any $w\in T_{q}M$ there exists a Jacobi field $Y$ along $\gamma$ such that $Y(0)=0$ and $Y(t)=w$. Notice that by our assumption on nonpositive radial curvature, we have ${\langle R(Y(t),\gamma^{{}^{\prime}}(t))\gamma^{{}^{\prime}}(t)},{Y(t)\rangle}\leq 0$ for any $t>0$ and $v\in S_{p}M$. We now recall the standard technique of Jacobi field comparison. Since the second derivative of the Jacobi field is related to the curvature tensors via the Jacobi equation, quantitative estimates are obtained by imposing conditions on curvatures. Note that
$\lambda(s)=\frac{\left\lVert Y(s)\right\rVert^{2}}{{\langle Y^{{}^{\prime}}(s)},{Y(s)\rangle}},\penalty 10000\ s>0$
is well defined since $M$ has no focal points. Using l’Hospital rule, we see that $\lambda(0)=0$ since $Y^{{}^{\prime}}(0)\neq 0$. Differentiating and using the Jacobi equation we have
$\lambda^{{}^{\prime}}(s)$ $=\frac{2{\langle Y^{{}^{\prime}}(s)},{Y(s)\rangle}^{2}-\left(\left\lVert Y^{{}^{\prime}}(s)\right\rVert^{2}-{\langle R(\gamma^{{}^{\prime}}(s),Y(s))\gamma^{{}^{\prime}}(s)},{Y(s)\rangle}\right)\left\lVert Y(s)\right\rVert^{2}}{{\langle Y^{{}^{\prime}}(s)},{Y(s)\rangle}^{2}}$
$\leq\frac{2{\langle Y^{{}^{\prime}}(s)},{Y(s)\rangle}^{2}-\left\lVert Y^{{}^{\prime}}(s)\right\rVert^{2}\left\lVert Y(s)\right\rVert^{2}}{{\langle Y^{{}^{\prime}}(s)},{Y(s)\rangle}^{2}}\quad\text{(nonpositivity of radial curvature)}$
$\leq\frac{2{\langle Y^{{}^{\prime}}(s)},{Y(s)\rangle}^{2}-{\langle Y^{{}^{\prime}}(s)},{Y(s)\rangle}^{2}}{{\langle Y^{{}^{\prime}}(s)},{Y(s)\rangle}^{2}}=1\quad\text{(Cauchy-Schwarz inequality)}$
Hence $\lambda(s)\leq s$ or equivalently,
$s{\langle Y^{{}^{\prime}}(s)},{Y(s)\rangle}\geq\left\lVert Y(s)\right\rVert^{2},\quad s>0$
The claim follows from (4). ∎
We are now ready to prove the main spectral result of this section. The bounds on the Laplacian and the bilaplacian demanded below are satisfied if lower bounds on curvatures and derivatives of metric tensor are assumed. It should be kept in mind that there exist Hadamard manifolds with pure point spectrum if the curvatures become unbounded.
Remark 3.7.
Note that we adopt the convention of considering the Laplacian as a positive operator, in contrast to the convention used in [Xav88] for Hadamard manifolds. Consequently, certain inequalities in the subsequent theorem appear with opposite signs when compared to the corresponding proof in [Xav88].
@@
Let $M$ be a simply connected Riemannian manifold without focal points. Suppose there exists a point $p$ in $M$ such that the radial curvatures with respect to $p$ are nonpositive. If there exist constants $c_{1},c_{2}>0$ such that $f$ and the Laplacian $\Delta$ satisfies
  1. (i)
    $\Delta f\geq-c_{1}$
  2. (ii)
    $\Delta^{2}f\leq c_{2}f^{-3}$
then the absolutely continuous spectrum of $\Delta$ is nontrivial. In fact, $(\alpha,\infty)$ is in the absolutely continuous part of the spectrum of $\Delta$ where $\alpha=\frac{3}{2}c_{1}+\frac{1}{4}c_{2}$
@@
Note that $\left\lVert\nabla f\right\rVert$ is bounded. Since $M$ has nonpositive radial curvature at $p$, from Lemma 3.6 we obtain that $\mathrm{Hess}\penalty 10000\ r^{2}(p)$ is the Riemannian metric at $p$ and for any $p\neq q\in M$,
$\mathrm{Hess}\penalty 10000\ r^{2}(q)(w,w)={\langle\nabla r(q)},{w\rangle}^{2}+r(q)\mathrm{Hess}\penalty 10000\ r(q)(w,w)\geq\left\lVert w\right\rVert^{2},\quad\forall\penalty 10000\ w\in T_{q}M,\left\lVert w\right\rVert=1$
From here, the arguments are standard, as in [Xav88, Theorem 1], for example. We write them down here for completeness. Recall that $f(q)=(1+r(q)^{2})^{1/2}$ and its Hessian can be computed to be
$\mathrm{Hess}\penalty 10000\ f(q)(w,w)$ $=-\frac{r^{2}(q)}{(1+r^{2}(q))^{3/2}}{\langle w},{\nabla r(q)\rangle}^{2}+\frac{1}{2(1+r^{2}(q))^{1/2}}\mathrm{Hess}\penalty 10000\ r^{2}(q)(w,w)$
$\geq\frac{1}{(1+r^{2}(q))^{3/2}}$
Taking $h(q)=\frac{1}{(1+r^{2}(q))^{3/2}}$, we see that
$\Delta h$ $=\frac{3}{f^{4}}\Delta f+\frac{12}{f^{5}}\left\lVert\nabla f\right\rVert^{2},$
$\implies\Delta h$ $\geq\frac{3\Delta f}{f^{4}}\geq-\frac{3c_{1}}{2}$
Thus for
$\beta=\inf\frac{1}{4h}\{2\Delta h-\Delta^{2}f\}\geq-\frac{3c_{1}}{2}-\frac{c_{2}}{4}$
and the claim now follows as in [Xav88, Theorem 3]. ∎
Clearly, the above theorem generalizes [Xav88, Theorem 1] under much weaker curvature conditions. However, more refined questions such as identifying whether the point spectrum or the singularly continuous spectrum occur are not amenable to this approach via the use of the distance function. To overcome this, we now turn our attention on the other natural geometrically defined convex function, the Busemann function.
4. A construction using Busemann functions and its applications
In this section, we construct a smooth strictly convex function out of Busemann functions on simply connected manifolds without focal points, using the geometry at infinity and assuming, in addition, strict negativity of Ricci curvature. As remarked in the introduction, this is motivated by a similar construction on symmetric spaces of noncompact type outlined in [DX06, penultimate paragraph p.657657]. The main purpose of this construction here is that it allows us to prove new results (Theorem 4.3 and Corollary 4.7) completely determining the nature of the spectrum for certain interesting classes of manifolds without focal points whose horospheres have constant mean curvature.